log(10x)^2=2+2logx This one is false right? Thank you for helping me.
Answer by venugopalramana(3286)

log(10x)^2=2+2logx This one is false right? Thank you for helping me.
LHS=2LOG(10X)=2{LOG(10)+LOG(X)}=2+2LOG(X)